OpenStudy (anonymous):

More algebra 2 help? I'm really bad at these word problems...

OpenStudy (anonymous):

The amount of radioactive isotope decays in half every 1 year. The amount of the isotope can be modeled by \[f(x)=346(\frac{ 1 }{ 2 })^x\] and f(1)=173. What does the \[\frac{ 1 }{ 2 }\] represent?

OpenStudy (anonymous):

A. The starting amount of the isotope B. The amount of the isotope after 1 year C. The rate the isotope decreases D. The number of years that have passed

OpenStudy (anonymous):

Very similar to my last question

OpenStudy (***[isuru]***):

hi, It's same as ur previous one .... It's represent the rate the isotope decays... it's C

OpenStudy (anonymous):

I thought so! I just needed reassurance :) thank you

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (anonymous):

Help with another one like this? I think I have the answer but I want to be sure I am right...

OpenStudy (anonymous):

The population of a species of rabbit triples every year. This can be modeled by \[f(x)=4(3)^x\] and \[f(5)=972\] What does the 5 represent? A. The starting population of the rabbits B. The population of the rabbits after 5 years C. The rat the population increases D. The number of years that have passed

OpenStudy (anonymous):

I think the answer is B... @bhowl77 ?

OpenStudy (***[isuru]***):

yes ... u r correct!!

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

Thankuuuuuuuu :DDDDDD